OpenStudy (anonymous):

Assume that x and y are both differentiable functions of t and find the required values of dy/dt and dx/dt. y = √x (a) Find dy/dt, given dx/dt = 3 and x = 4. dy/dt = (b) Find dx/dt, given x = 25 and dy/dt = 2. dx/dt =

OpenStudy (anonymous):

It means derivative of y, right?

OpenStudy (anonymous):

Ok, so when I find the derivative I get 1/2x^-1/2... but I don't know what to do next

OpenStudy (dumbcow):

its the chain rule \[\frac{dy}{dt} = \frac{dy}{dx}*\frac{dx}{dt}\]

OpenStudy (dumbcow):

\[= \frac{1}{2\sqrt{4}}*3\]

OpenStudy (anonymous):

Ok. I understand that. You plugged 4 into the derivative. Then we multiply and get \[3/2\sqrt{4}\] right? And that's the answer?

OpenStudy (dumbcow):

yes well simplify it to 3/4

OpenStudy (anonymous):

Ok. Wow, that's not that hard. Thanks.

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

For b my book says the answer should be 20 but I'm getting 1/10.

OpenStudy (dumbcow):

\[2 = \frac{1}{2\sqrt{25}}*\frac{dx}{dt}\]

OpenStudy (dumbcow):

see why it is 20

OpenStudy (anonymous):

Oops thanks.

OpenStudy (dumbcow):

dont switch the dx and dy , always start with \[\frac{dy}{dt} = \frac{dy}{dx}*\frac{dx}{dt}\] then substitute given values

OpenStudy (anonymous):

Yeah, I understand. I got dx/dt by itself but I lost the two somewhere so I didn't divide by it.

OpenStudy (anonymous):

Thank you. =)